Hawaii-2017-01 experiment with new MOBYrefresh blue Resonon - BS01cfg006 and 007

Intial testing of BS01 for deployment in MOBY262. On day 1 Iris blocks are added to Track1, 9, 11 and 13 to reduce the signal and increase integration times for in-water data collection (cfg006). On 7 Feb Mark dis-assembled the secured the 4x iris settings, and re-assembled the BS01 (cfg007). Day 2 Mike retook some data to confirm the iris-ed tracks throughput. Day 3 was a repeat of the day 1 Lu data but with the heads at the correct distance (a real cal). And day 4 was the Es head cal with a repeatablilty measurement when you have to disconnected the end-block and re-attached said end-block.

Problems seen (and solved) and good things:
  1. UV ghosting found from track 1 onto 14. Page Number 5.06.

Log Sheets

Photos

Es responses
Lu responses

All the raw data files and their images , Table of KEYWORDS variable
Page Number
Link
Description
Date
1.01
Day 01 - 6 Feb Raw Data , Table
DAY01 - Oriel, 4x iris @ shutter block adjustment for Tracks 1,9 11 and 13 - cfg006
Feb 6, 2016
1.02
Day 02 - 8 Feb Raw Data , Table
DAY02 - Oriel, 4x iris @ shutter block confirming adjustment for Tracks 1,9 11 and 13 (after Mark secured iris settings) - cfg007
Feb 9, 2016
1.03
Day 03 - 10 Feb Raw Data , Table
DAY03 - OL425 lamp data of LuTop, Mid and Bot with heads, shutters and iris - cfg007
Feb 10, 2016
1.04
Day 04 - 11 Feb Raw Data , Table
DAY04 - Lamp cals of Es, LuTop, Mid and Bot with heads, shutters and iris - cfg007
Feb 10, 2016
 
Day01 - Oriel, 4x iris @ shutter block adjustment for Tracks 1,9 11 and 13 - cfg006
2.01 iris block Confirming Mikes Iris Block numbers Feb 8, 2017
2.02 finding tracks Looking at track definitions Feb 8, 2017
2.03 Saddle shape changes Changes to saddle shape with iris addition Feb 8, 2017
Day02- Oriel, 4x iris @ shutter block confirming adjustment for Tracks 1,9 11 and 13 (after Mark secured iris settings) - cfg007
3.01 iris block check Confirming Mikes Iris Block numbers after Mark secured iris settings Feb 9, 2017
Day03 - OL425 lamp data of LuTop, Mid and Bot with heads, shutters and iris - cfg007
4.01 rough rsp A rough intial system response of the BS01 with iris, shutters and Lu heads Feb 10, 2017
Day04 - Lamp cals of Es, LuTop, Mid and Bot with heads, shutters and iris - cfg007
5.01 Es rsp Es rsp before the disconnected end-block Feb 11, 2017
5.02 Es rsp repeatability Es rsp before and after the disconnected end-block Feb 11, 2017
5.03 Lu rsp Lu rsp Feb 11, 2017
5.04 int time cal My intial look at Mike int cal data Feb 11, 2017
5.05 %std Es and Lu Looking the %std of Es repeatability for Es and Lu net signals Feb 13, 2017
5.06 Ghosting check Checking to see if light from track 1 is showing up on track 14 in the UV 19 Jan 2018
       

 

DAY 1 EMAIL from Mike

Hi Steph,
Well, here we go!
Mark/Sean got the shutter block together, with 4x iris to adjust the sensitivity of Es and 3x Lu collectors for the MOBY262 deployment.  So, I collected some *.fits sig & bac data today, via PC & Andor SOLIS software. There is 1x log sheet and 2x pages of my notes @ /ftp1/Mike/HI-2017-01/BS01data/doc,
plus photos at .../photos/ + *.fits data at .../BS01data/day01/.

The 4x ports with iris are #1, 9, 11, 13.  #1 will get Es, #9, 11, 13 = Lu. In the photo #06,7,8 on can see the fiber-bundle / collumator / shutter for track #1.  The iris is the knurled (not smooth) barrel near the shutter. The neighbor trk #2 & 10 can be seen to not have an iris...
Photos #9,10,11 try to show the iris @ port 9,11,13.

Adding the iris, even with the iris fully open, attenuated the light signal by ~2.5 x by my reckoning.  This can be seen in s,b*01.fits, by comparing the signal for non-iris tracks 2:8,10,12,14 vs signal for iris trks 1,9,11,13.

Mark said that even with the iris fully open, there is some iris in the FOV. Could you check my estimate of 2.5x decreased sensitivity for trk 1,9,11,13? I did this by approximating center-of-saddle at near max sig ADU. I think I did this via sig scans, but should have done this on net-signal ADU...

Mark asked what the maximum iris attenuation is. So, fits 02 = only shtr 1,9,11,13 = open, 4sec exp, all iris open, and fits 03 = all iris closed.  My guess was a ~325x net sig attenuation. Could you please check this?

sig,bac #04 was only trk #1 shtr open. Then I adjusted the iris to get an additional 40x net-signal decrease - my thinking was that the open iris gave a 2.5x decrease, so another 40x should give a total attenuation of 100x for Es.  I know the doc we all worked up said that
Es should have a 200x decrease to get 60 sec int time in-water, but 200x just seemed TOO MUCH to me!

Here's where I messed up: I did not save s,b scan #5 after I adjusted the trk #1 iris - I think scan #09, later should do though...

Next, scan #5 was only shtr 9,11,13 = open.  Then I tried to knock down the net sig by another 4x for Lu, for a total of 10x decrease (2.5x from open iris * 4x from adjusted iris = 10x total attenuation). Scan #6 = after adj iris 9, #7 after adj iris 11, #8 after adj iris 13.

Then, scan #9 = shtr open = trk 1,9,11,13, iris adj = 1,9,11,13, at 4 sec. Could you please compare scn #9 vs scn #2 for iris attenuation factors?

Scan #10 = all shtr open @ 1 sec, to compare with scan #01. Hopefully trk #1 net sig is down 100x, and trk 9,11,13 are down 10x, compared with an average of other tracks without iris... Could you please check this?

Scan #11 = lamp off = ambient, shows the room fluor light leak. I could do a better job of sealing the sphere and black-cloth-ing the fiber bundle if we need to repeat this.  There was no black cloth used today so as to ease my access to the iris's...

If this iris adjustment seems ok, then Mark will need to dis-assemble the shutter block to lock down the iris positions.  Then all will need re-assembled, and re-checked.

Luck to you!  Thanks to you!  MF

DAY 2 EMAIL from Mike

Hi Steph,
Mark dis-assembled the BS01cfg006 last night, secured the 4x iris settings,
and re-assembled the BS01 today ... as BS01cfg007,
then I SOLIS-collected some data to re-look the iris-ed tracks throughput.

On your FTP site I re-named yesterday's photos, and added a pair from
Art's test earlier this month, and added photos from last night of the
torn-apart BS, and added photos from today of the data acq. setup.
And, I re-scanned my updated log sheet(s) from yesterday and today.

The one setup change today vs yesterday was I added a black cloth
around the shutter block to try to keep out ambient room light.

So, today: scan set #1, s,b2017020901.fits was all shtr open, 1sec;
scn 02 was only shtr 1,9,11,13 open, 1sec;
scn 03 was only shtr 1,9,11,13 open, 8sec
(NOTE: the 8 sec was double what it was yesterday to get ~full scale ADU)
scn 04 was all shtr closed, 8 sec, to get an approximate "dark"
scn 05 was only shtr 1 open, 60 sec, to get a saddle shape at ~full scale ADU
scn 06 was lamp off, all shtr open, 60 sec, to look at ambient light leak.

Luck!  MF

DAY 3 EMAIL from Mike

Hi Steph,
There are some more .fits files at /ftp1/Mike/HI-2017-01/BS01data/day03/ !
For each of the three Lu's, there were 5x pairs of sig&bac scans,
with only one shutter-block shutter open for each Lu, where
LuTop = shtr/trk #9, Mid = #11, Bot = #13.
I also collected 3x darks = d2017-211xx.fits for each Lu, which was the
all shutter-block shutters closed, and a sig or lite scan via the camera
while the lamp was still on - to see if the shutter-block shutters were leaking light...

There is one log sheet to go with this.

I have lamp Voltage files, but I'll look at those tomorrow,
and I have some photos, which I'll look at and rename now
then give you an email when they are FTP'ed.
Luck!  Thanks!  MF

DAY 3 EMAIL from Mike part 2

Hi Steph,
The OL425 has a photodiode monitor, and the Lamp#9 NIST cal radiance you have for it were at PD mon = 26.33 fL (footLamberts). Tonight's PD readings were ~26.39 fL, but I wouldn't worry aboot trying to normalize for this because: 1. it is schmall (~0.2%), and 2. I typically set the OL425 at 6 inch distance from a Lu head so as to not load the sphere with reflections off the Lu window, but tonight I used a jig between OL425 & Lu so the PD readings are probably not much use...
MF

DAY 3 EMAIL from Mike part 3

Ahoy,
The first BS Lu pseudo-cal went well and Stephanie has the data. I'd say MOBY is ready to tape 'n paint...

I pulled most of the slack out of the BS Es fiber hose, so that might be ready to snug down above the water line, and maybe the FO channel is OK to button up?

One curious ting aboot the Lu cals: Top needed a 20 sec exposure time, Mid needed 15 sec, and Bot needed 6 sec - to get similar signal ADU. Maybe this is a function of the iris settings, or maybe it is a function of the lengths of the new 600 um fiber, or some combo?  Good news is that exposures are much above the 35 milisec shutter open/close times...
MF

DAY 4 EMAIL from Mike

Hi Steph,
There are some more data that need your help! Tonight should be the last night of BS data. Mark plans to finish BS assembly & get the housing on tomorrow. Once the housing is on only Art has access to BS control...

Tonight I got an Es cal via FEL-F600(Cal#3) which should be the cal irradiance used w/ pre & post-MOBY260.  I'll probably need to send you a good BBfit file, but the data at M260 waves should be good for now.

There were 5x sig,bac pairs of FEL scans, then I disconnected the end-block - where the bottom of the BS fibers terminate - then re-attached said end-block, and collected another 5x sig,bac pairs. The idea was to see if the end-block attachment is repeatable, because that is what will happen tomorrow.

Then, I repeated the Lu cals via the OL425(L9) - like yesterday, except at the proper 6 inch distance between the sphere and Lu head.  The PD monitor fL values should be used to normalize, but that is probably not extremely necessary right now. The fL value during NIST cal to normalize to was 26.33 fL. All tonight's data were 26.05 to 26.12 fL, so tonight's net signals should be multiplied by a factor of 26.05/26.33 = 0.9894, for example. The PD fL's are on the log sheet.

Last, I did a quick exposure time cal.  This processing is likely not too obvious, and I'll not be able to explain it right now, so this can wait till tomorrow!

This all makes for a total of 52x sig,bac pairs of .fits files. There are 2x log sheets. I have some photos, but those will also have to wait until tomorrow.
Thanks, MF

EMAIL from Mike on BS01 calibrations and lamp files

Hi STeph,
Here are the ASSkey text files for the pre-M262 BS01 responses:
<attached: F600c3.txt, OL425L9.txt>
Both have the NIST FASCAL spectra at the course NIST wavelength scale
and the Black-Body fit at 0.25 nm interval - you'll want to use the BBFits.

For the Lu cals, the PhotoDiode monitor on the OL425 @ pre-M262 was aboot 26.1 fL,
(i.e. see the "fL" column in attached BS01cfg006_log_05.jpg)
but the PD mon @ NIST FASCAL was 26.33 fL, so all the pre-M262 BS ADU
need adjusted by aboot 26.33/26.1 = 1.0088 before calculating the response -
ex. Lt_scan#1_rel = Lt_scan#1_ADU * 26.33/26.09,
Lt_scan#1_per = per(Lt_scan#1_rel, Lt_exp_time),
Lt_scan#1_resp = interpolated_BBFit_radiance / Lt_scan#1_per
(but you knew that).

MF

DAY 4 EMAIL from Mike on int time cal
Hi Steph,
I looked at my processing for the last 2 times doing Int Time Cals for pre & post M260 LuBot,
and, I'm not so sure that my approach is an optimum approach,
so you might want to look at what I have to say below and try to improve upon it!

First, "let it be know that" I collected BS data at 3x OL455 sphere settings,
which I call "Brite", "Mid", and "Dim".  The shortest exp times were via Brite lamp,
and longest exp times were at Dim lamp.

Second, it wanted to see if what I was aboot to tell you worked, so I ran it through myself.
I processed the Net Signal = Sig minus Bac, in ADU,
using the mean for LuBot Trk#13 = pix 893:936 (your webpage H17-01, page 2.02 "finding tracks").
I used file s,b2017021226...52.fits, which gave 27x net's.

Here is my .m scan-navigation cheat-sheet:  (this lines up better in MATLAB editor...)

% scn  1   2   3   4  5  6   7  8  9 10  11 12 13 14  15 16 17 18 19 20 21 22  23 24 25 26 27
% sec  1  .05 .07 .1 .2  1  .3 .5 .7 1   1  2  3  1   5  7  1  10 10 20 30 10  50 10 70 10 100
% BMD  B   B   B   B  B  B   B  B  B B   M  M  M  M   M  M  M  M D  D  D  D   D  D  D  D  D
% 1sec *                 *           *   *        * *

Here is my loader for the PD monitor data:

% OL455 PD monitor (Amps)
%-------------------------
%     1     2     3     4     5     6     7     8     9     10
pd=([.8521 .8524 .8521 .8522 .8524 .8522 .8522 .8523 .8521 .8522]+1)*1e-6  , whos pd  %  [1x10]
%         11    12    13    14    15    16    17    18
pd=[pd ([.8862 .8857 .8856 .8856 .8853 .8853 .8851 .8852]+1) *1e-7 ], whos pd  %  [1x18]
%         19    20    21    22    23    24    25    26    27
pd=[pd ([.8829 .8828 .8828 .8825 .8822 .8822 .8823 .8823 .8823]+1) *1e-8 ], whos pd  %  [1x27]

NOTE: I did not end up using the PD data, but I think the processing could be done
to make use of these, because the transitions from Brite to Mid, and Mid to Dim
agree well between the BS Net Sig ratios and the PD mon ratios
<attached: rb62bs01_exp_1.png>

I checked the Net Sig repeats at 1 sec for the Brite lamp (+/- 0.5%),
and at 1 sec for the Mid lamp (+/- 1%),
and at 10 sec for the Dim lamp (+/- 1.5%).
Since they seemed OK, and the PD mon repeats were essentially 1,
I did my Net Sig "normalizing" to the nearest-in-time
1 sec exposure Net Sig at Brite & Mid lamp settings,
and to the nearest-in-time 10 sec net at Dim lamp setting
(hopefully this will make sense below!).

Here is what I used for Nominal Exposure Time:

% N      1   2  3  4  5  6  7   8   9 10 11 12 13 14 15 16 17 18
nom = [.05 .07 .1 .2 .3 .5 .7   1   2  3  5  7 10 20 30 50 70 100], whos nom  %  [1x18]
% scn=[  2   3  4  5  7  8  9  10  12 13 15 16 18 20 21 23 25 27 ];

So, there will be 18x "Calibrated" exposure times for these 18x nominal times.
I get Calibrated times as "rel", by normalizing to 1 sec net signal:

% Brite
rel(:,01)=net(:,02) ./ net(:,01);  % #02=  0.05 / 1 sec
rel(:,02)=net(:,03) ./ net(:,01);  %  03=  0.07
rel(:,03)=net(:,04) ./ net(:,06);  %  04=  0.1
rel(:,04)=net(:,05) ./ net(:,06);  %  05=  0.2
rel(:,05)=net(:,07) ./ net(:,06);  %  07=  0.3
rel(:,06)=net(:,08) ./ net(:,06);  %  08=  0.5
rel(:,07)=net(:,09) ./ net(:,10);  %  09=  0.7
rel(:,08)=net(:,10) ./ net(:,10);  %  10=  1
%
% Mid
rel(:,09)=net(:,12) ./ net(:,11);  % #12=  2 / 1 sec
rel(:,10)=net(:,13) ./ net(:,14);  %  13=  3
rel(:,11)=net(:,15) ./ net(:,14);  %  15=  5
rel(:,12)=net(:,16) ./ net(:,17);  %  16=  7
rel(:,13)=net(:,18) ./ net(:,17);  %  18= 10
%
% Dim                                Mid,10sec    Mid,1sec
rel(:,14)=net(:,20) ./ net(:,19) .* (net(:,18) ./ net(:,17));  % #20=  20 / 10
rel(:,15)=net(:,21) ./ net(:,22) .* (net(:,18) ./ net(:,17));  % #21=  30
rel(:,16)=net(:,23) ./ net(:,24) .* (net(:,18) ./ net(:,17));  % #23=  50
rel(:,17)=net(:,25) ./ net(:,26) .* (net(:,18) ./ net(:,17));  % #25=  70
rel(:,18)=net(:,27) ./ net(:,26) .* (net(:,18) ./ net(:,17));  % #27= 100

For the Dim scans, "one can see that" (I hate when they say that!)
I added a correction to get from 10 sec at Mid to 1 sec at Mid (net#18/net#17) -
this is where I think there might be a better way of doing this -
maybe by using the PD monitor to correct between the Brite & Mid & Dim settings ?

Plotting the "rel" net signals on a semilogy it looked like pix 300:900
was a stable range, so I got the mean and std over pix 300:900:

N=[300:900]; nanmean(N),clear ans  %  600
for i=1:size(rel,2)
int(i)=nanmean(rel(N,i));  std(i)=nanstd(rel(N,i));
end, clear i, whos int* std*  %  [1x18]

[[1:18]; nom;  int; std./int*100]'
%  N    nom      avg        %std
% --    ----    --------   -------
%  1      0.05   0.094278  1.0278
%  2      0.07   0.11343   0.85473
%  3      0.1    0.14175   0.78595
%  4      0.2    0.23689   0.53124
%  5      0.3    0.33199   0.37709
%  6      0.5    0.52217   0.29506
%  7      0.7    0.71216   0.23327
%  8      1      1         0
%  9      2      1.9524    0.63782
% 10      3      2.9134    0.55747
% 11      5      4.8333    0.55972
% 12      7      6.7566    0.62245
% 13     10      9.6539    0.66262
% 14     20     19.24      1.1847
% 15     30     28.909     0.97703
% 16     50     48.403     0.85737
% 17     70     67.533     1.1422
% 18    100     96.7       1.2415

Which plotted up like this:
<attached: rb62bs01_exp_2.png>

I do NOT think this std is a valid uncertainty estimate,
because it does not include the unc for the transition
between 10 sec to 1 sec at Mid lamp setting...

Here are my results:
<attached: BS01_exptimecal(rev13Feb2017).txt>

I think we can use this to get started, and worry aboot improving
upon this approach when we have some more time...
MF